# From System 1 to System 2: A Survey of Reasoning Large Language Models

### v2

@@ 3.1.1 Output Behaviour Perspective
Longer Inference Length & Time: Recent research [49, 50, 51, 52, 114] indicates that reasoning LLMs often generate outputs exceeding 2000 tokens to tackle complex problems in coding and mathematics. However, this extended output length can sometimes lead to overthinking, where the model spends excessive time on a problem without necessarily improving the solution. Studies [49] highlight that while autoregressive generation and Classic CoT can effectively solve simpler problems, they struggle with more complex tasks. Research [115, 116] shows that in multimodal domains, many problems demand careful observation, comparison, and deliberation. Additionally, Search-o1 [117] suggests that slow-thinking mechanisms are particularly beneficial in areas requiring external knowledge or where potential knowledge conflicts arise. In medical scenarios, complex problems, such as those requiring test-time scaling techniques, demonstrate significant improvements [52].
@@ 3.2 Core Method
TABLE I: Summary of Structure Search method based on the definition of actions and rewards.
Category Reasoning LLMs Characteristic
Actions Reasoning Steps as Nodes RAP [14], ORM [124], Forest-of-Thought [125] Actions represent intermediate reasoning steps.
Token-level Decisions CodeTree [126], SPaR [127], TreeBoN [128] Actions involve generating tokens.
Task-specific Structures CWM [129], LLM-MCTS [130] Actions are domain-specific.
Correction and Exploration RethinkMCTS [131], MCTSr [132] Actions focus on revisiting and refining previous steps.
Rewards Outcome-based Rewards MC-NEST [133] Correctness or validity of the final outcome.
Stepwise Evaluations RAP [14], SRA-MCTS [134] Rewards are assigned at intermediate steps.
Self-evaluation Mechanisms SPaR [127], TreeBoN [128], MindStar [135] Rewards rely on the model’s own confidence.
Domain-specific Criteria LLM-MCTS [130], SR-MCTS [136] Rewards are tailored to specific tasks.
Iterative Preference Learning LLaMA-Berry [137], Marco-o1 [112], ReST-MCTS* [138] Rewards derive from comparing multiple solutions.
@@ 3.2.1 Structure Search
MCTS, a powerful search and optimization algorithm, effectively addresses these challenges by providing a structured framework to explore and evaluate reasoning paths systematically. It operates by constructing a reasoning tree, where each node represents a reasoning state, and actions expand the tree by considering potential next steps. Through the simulation of future states and the iterative backpropagation of estimated rewards, MCTS allows foundational LLMs to efficiently identify high-reward reasoning paths, mirroring human planning processes. This approach aligns with the core principles of reasoning LLMs, where thorough analysis and deliberate exploration are essential for generating well-reasoned outputs. Recent methods, such as RAP [14], enhance foundational LLMs by integrating MCTS with a world model, enabling the system to iteratively refine intermediate reasoning steps and improve future predictions. Similarly, Forest-of-Thought [125] utilizes MCTS to dynamically explore multiple reasoning trajectories, revisiting flawed paths and refining outcomes.
@@
The integration of MCTS into LLMs depends on defining actions and rewards to guide reasoning path exploration and assess quality. As shown in Table I, we classify the actions in prior work into four categories:
@@
Additionally, as illustrated in Table I, we classify the reward design into five categories:
@@
ORM DIVERSE[146] Prompting Fine-tuning SFT Multiple Reasoning Tasks Weighted Voting Verifier
@@
OVM[150] Sampling Feedback-guided SFT Math Reasoning Guided Decoding
@@
PRM ORPS[155] Sampling Feedback-guided SFT Code Generation Supervising Outcome Refinement
@@ 3.2.2 Reward Modeling
Two primary training paradigms are used to tackle multi-step reasoning tasks: outcome supervision and process supervision. Outcome supervision emphasizes the correctness of the final answer at a higher level of granularity, and the resulting model is referred to as the Outcome Reward Model (ORM) [32, 158]. In contrast, process supervision provides step-by-step labels for the solution trajectory, evaluating the quality of each reasoning step. The resulting model is known as the Process Reward Model (PRM) [159, 37, 160]. The main distinction between ORM and PRM is illustrated in Figure 5.
Figure 5: The comparison between ORM and PRM for assessing a complete solution trajectory. ORM only provides a single reward based on the correctness of the final answer, while PRM evaluates the quality of each reasoning step throughout the process.
PRM offers significant advantages [161, 147] in complex reasoning tasks for several key reasons. First, it provides fine-grained, step-wise supervision, allowing for the identification of specific errors within a solution path. This feature is especially valuable for RL and automated error correction. Second, PRM closely mirrors human reasoning behavior, which relies on accurate intermediate steps to reach correct conclusions. Unlike ORM, PRM avoids situations where incorrect reasoning can still lead to a correct final answer, thus ensuring more robust and interpretable reasoning. While PRM has primarily been applied to complex mathematical problems, its benefits have recently driven applications in other fields. For instance, ORPS [155] utilizes PRM to address complex code generation challenges, while Step-DPO [156] combines process supervision with the Direct Preference Optimization (DPO) algorithm [162] to improve long-chain mathematical reasoning. A summary of Reward Modeling method is presented in Table II.
@@ 3.2.3 Self Improvement
Reasoning LLMs exemplify a progression from weak to strong supervision, while traditional CoT fine-tuning faces challenges in scaling effectively. Self improvement, using the model’s exploration capabilities for self-supervision, gradually enhances LLMs performance in tasks such as translation [167], mathematics [163, 168], and multimodal perception [171]. This approach fosters exploration and application within reasoning LLMs [184, 185, 186, 166]. A summary of Self Improvement method is presented in Table III.
@@ 3.2.5 Reinforcement Fine-Tuning
Methods Model Attribute Incentivize Attibute Application & Benchmark
@@
DeepSeek-R1-Zero[31] DeepSeek-V3 TT Rule-Outome-Reward GPRO RL 800K Multiple Tasks
@@
Kimi v1.5[215] II TT Rule-Outcome-Reward $\text{PPO}^{*}$ RL & SFT Multiple Tasks
@@
RLHF-V[230] OmniLMM-12B II TT Process-Reward DDPO RL 1.4K Multiple Tasks
RLAIF[231] PaLM 2 Extra-Small TT Rule-Outome-Reward RLAIF RL Summary and Conversation Generation
@@
R1V[234] Qwen2-VL,Qwen2.5-VL II TT Rule-Outome-Reward GRPO RL 70K/70K/8K Multiple Tasks
VLM-R1[235] Qwen2.5-VL II TT Rule-Outome-Reward GRPO RL 120K Multiple Tasks
LMM-R1[236] Qwen2.5-VL II TT Rule-Outome-Reward PPO/RLOO RL 8K Multiple Tasks
Open-R1-Video[237] Qwen2-VL-7B II TT VV Rule-Outome-Reward GRPO RL 4K Multiple Tasks
Easy-R1 [238] Qwen2.5-VL II TT Rule-Outome-Reward GRPO RL 3K Multiple Tasks
@@
Demystify-LongCoT[239] Llama-3.1-8B, Qwen2.5 -7B-Math TT Rule-Outcome-Reward PPO/Reinforce++ RL & SFT 7.5K Multiple Math, MMLU
@@
MD-CoT[241] Process-Reward PPO RL
Despite its strengths, RFT faces the following challenges:
@@
    Unstable Long-CoT Generation: Long reasoning chains generated by RFT are prone to instability, including context overflow, failure to return final answers, and sensitivity to reward shaping [122]. For instance, methods like [239] inadvertently introduce cosine reward functions, which degrade performance with increased iterations. O1-Prune [244] uses post-hoc length pruning techniques [215] (via RL/SFT) to stabilize outputs.
@@
    Theoretical and Empirical Analysis: It is essential to clarify the relationship between RL training and the capabilities of the base model. For instance, it should be determined whether emergent properties (e.g., Long-CoT) arise from RL optimization or are latent traits of the base model. Systematic studies on reward design principles (e.g., sparse vs. dense rewards, multi-objective balancing) should be conducted to avoid unintended behaviors such as reward hacking.

### v3

@@ 3.1.1 Output Behaviour Perspective
Longer Inference Length & Time: Recent research [49, 50, 51, 52, 114, 115] indicates that reasoning LLMs often generate outputs exceeding 2000 tokens to tackle complex problems in coding and mathematics. However, this extended output length can sometimes lead to overthinking, where the model spends excessive time on a problem without necessarily improving the solution. Studies [49] highlight that while autoregressive generation and Classic CoT can effectively solve simpler problems, they struggle with more complex tasks. Research [116, 117] shows that in multimodal domains, many problems demand careful observation, comparison, and deliberation. Additionally, Search-o1 [118] suggests that slow-thinking mechanisms are particularly beneficial in areas requiring external knowledge or where potential knowledge conflicts arise. In medical scenarios [119], complex problems, such as those requiring test-time scaling techniques, demonstrate significant improvements [52].
@@ 3.2 Core Method
TABLE I: Summary of tree search-based methods. Tag:  N   =Natural language,  C   =Code,  E   =Math Expression, or  A   =Action.
Node
Method Format Partial Evaluate Rollout Tasks LLM Models
RAP [14] NN LLM Self-correction LLM-based Prediction
Planning, Reasoning
LLaMA-33B
ORM [126] NN Value/Reward Function N/A
Multiple Tasks
LLaMA-2-7B, GPT-2-small
Forest-of-Thought [127] NN LLM Self-correction Self-refinement & Iterative Improvement
Planning, Reasoning
LLaMA-3, Mistral7B, GLM-4-9B
CodeTree [128] NN   CC Execution Accuracy + LLM Self-correction Code Execution Code Generation GPT-4
TreeBoN [129] NN Value/Reward Function Speculative & Dynamic Strategies
Planning, Reasoning
LLaMA-3-8B
CWM [130] CC Compare with Golden Data Code Execution Alignment Task N/A
LLM-MCTS [131] NN LLM Self-correction + Policy LLM-Based Prediction Household Environments
GPT-2, GPT-3.5
RethinkMCTS [132] CC Execution Accuracy Code Execution Code Generation
GPT-3.5-turbo, GPT-4o-mini
MCTSr [133] NN LLM Self-correction Self-Refinement & Iterative Improvement Mathematical Reasoning LLaMA-3-8B
MC-NEST [134] NN LLM Self-correction Reasoning Path Generation Mathematical Reasoning
GPT-4o, Phi-3-mini
SRA-MCTS [135] NN Execution Accuracy Reasoning Path Generation Code Generation LLaMA-3-70B-Instruct
SPaR [136] NN LLM Self-correction Self-Refinement & Iterative Improvement Instruction Following LLaMA-3-8B
MindStar [137] NN Value/Reward Function Reasoning Path Generation Mathematical Reasoning
LLaMA-2-13B, Mistral-7B
SR-MCTS [138] EE Compare with Golden Data Math Expression Generation Financial Fraud Detection GPT
LLaMA-Berry [139] NN Compare with Other Solutions Math Expression Generation Mathematical Reasoning LLaMA-3.1-8B
Macro-o1 [112] NN   AA LLM’s Output Probabilities Reasoning Path Generation
Multiple Tasks
Qwen2-7B-Instruct
ReST-MCTS* [140] NN Probability to Correct Answer Reasoning Path Generation Mathematical Reasoning
LLaMA-3, Mistral-7B, SciGLM-6B
CoMCTS [141] NN Compare with Other Solutions Reasoning Path Generation Multiple Tasks
GPT-4o, Qwen2-VL-7B, LLaMA-3.2-11B
C-MCTS [142] NN Compare with Golden Data Math Expression Generation Mathematical Reasoning Qwen2.5
rStar-Math [143] EE Compare with Other Solutions Math Expression Generation Mathematical Reasoning
Phi3-mini-Instruct, Qwen2.5-Math-1.5B, Qwen2.5-Math-7B
AStar[144] NN Compare with Other Solutions Reasoning Path Generation Multiple Tasks
Qwen2.5-7B, Qwen2-VL-2B, Qwen2-VL-7B
DeepSolution [145] NN   AA Compare with Golden Data Reasoning Path Generation Multiple Tasks Qwen2.5-7B-Instruct
VisuoThink[146] NN Compare with Golden Data Reasoning Path Generation Multiple Tasks
GPT-4o, Qwen2-VL-72B-Instruct, Claude-3.5-sonnet
TongGeometry [147] EE Compare with Golden Data Math Expression Generation Mathematical Reasoning GPT-4
PPO-MCTS [148] NN Compare with Other Solutions Reasoning Path Generation Alignment Task LLaMA-7B
@@ 3.2.1 Structure Search
MCTS, a powerful search and optimization algorithm, effectively addresses these challenges by providing a structured framework to explore and evaluate reasoning paths systematically. It operates by constructing a reasoning tree, where each node represents a reasoning state, and actions expand the tree by considering potential next steps. Through the simulation of future states and the iterative backpropagation of estimated rewards, MCTS allows foundational LLMs to efficiently identify high-reward reasoning paths, mirroring human planning processes. This approach aligns with the core principles of reasoning LLMs, where thorough analysis and deliberate exploration are essential for generating well-reasoned outputs. Recent methods, such as RAP [14], enhance foundational LLMs by integrating MCTS with a world model, enabling the system to iteratively refine intermediate reasoning steps and improve future predictions. Similarly, Forest-of-Thought [127] utilizes MCTS to dynamically explore multiple reasoning trajectories, revisiting flawed paths and refining outcomes. We list the recent methods in Table I.
@@
The integration of MCTS into LLMs depends on defining actions and rewards to guide reasoning path exploration and assess quality. We classify the actions in prior work into four categories:
@@
Additionally, we classify the reward design into five categories:
@@
ORM Verifier Training[32] Existing Data Verification SL Math Reasoning GSM8K, ORM
ORM PRM Comparison[156] Human Annotation Feedback-guided SFT & RL Math Reasoning PRM ORM Analysis
OVM[157] Sampling Feedback-guided SFT Math Reasoning Guided Decoding
ENCORE[158] Existing Data Entropy-guided Training-free Safety Tasks Entropy of Safety Attribute
PRM Outcome-based DIVERSE[159] Prompting Fine-tuning SFT Multiple Reasoning Tasks Weighted Voting Verifier
@@
ORPS[163] Sampling Feedback-guided SFT Code Generation Supervising Outcome Refinement
R-PRM[164] Sampling Feedback-guided SFT & RL Math Reasoning Reasoning-Driven Supervision
BiRM[165] Prompting Feedback-guided SFT Math Reasoning Bidirectional Reward Signals
DeepSeek-GRM[166] Sampling Feedback-guided SFT & RL Multiple Reasoning Tasks Inference-Time Scalability
RewardAgent[167] Existing Data Feedback-guided SFT & RL NLP Tasks Human & Verifiable Signals
PAR [168] Sampling Feedback-guided SFT & RL NLP Tasks Centered Reward Shaping
SCIR [169] Sampling Feedback-guided SFT & RL NLP Tasks Self-Consistency Enforcement
@@
Multimodal MSTaR[174] Sampling Self-training SFT Multiple Reasoning Tasks Adaptive Temperature Adjustment
VisualPRM[175] Sampling Fine-tuning SFT Math Reasoning Multimoda PRM, BoN
UnifiedReward[176] Sampling Feedback-guided SFT & RL Multimodal Tasks Unified Multimodal Reward Modeling
Others Pro-Out Feedback[177] Existing Data & Annotation Feedback-guided SFT & RL Math Reasoning Process & Outcome Supervision
Verify Step-by-Step[178] Human Annotation Feedback-guided SFT Math Reasoning Process Reward Annotation
@@ 3.2.2 Reward Modeling
Two primary training paradigms are used to tackle multi-step reasoning tasks: outcome supervision and process supervision. Outcome supervision emphasizes the correctness of the final answer at a higher level of granularity, and the resulting model is referred to as the Outcome Reward Model (ORM) [32, 181, 182]. In contrast, process supervision provides step-by-step labels for the solution trajectory, evaluating the quality of each reasoning step. The resulting model is known as the Process Reward Model (PRM) [156, 37, 183].
PRM offers significant advantages [184, 160] in complex reasoning tasks for several key reasons. First, it provides fine-grained, step-wise supervision, allowing for the identification of specific errors within a solution path. This feature is especially valuable for RL and automated error correction. Second, PRM closely mirrors human reasoning behavior, which relies on accurate intermediate steps to reach correct conclusions. Unlike ORM, PRM avoids situations where incorrect reasoning can still lead to a correct final answer, thus ensuring more robust and interpretable reasoning. While PRM has primarily been applied to complex mathematical problems, its benefits have recently driven applications in other fields. For instance, ORPS [163] utilizes PRM to address complex code generation challenges, while Step-DPO [179] combines process supervision with the Direct Preference Optimization (DPO) algorithm [185] to improve long-chain mathematical reasoning. The framework has also shown promise in multimodal settings, with M-STAR [174] demonstrating how self-evolving training can optimize PRM for vision-language tasks, and VisualPRM [175] establishing effective process reward modeling for multimodal reasoning through its Best-of-N evaluation approach. A summary of Reward Modeling method is presented in Table II.
@@
PSRLM [196] Sampling Language Model RL Reasoning
SCRIT [197] Sampling Language Model SFT Arithmetic Reasoning
S²R [198] Sampling Language Model SFT & RL Arithmetic Reasoning
Self-Training [199] Sampling Language Model SFT Arithmetic Reasoning
STL [200] Sampling Language Model SFT & RL State-Value Estimation
Genius [201] Sampling Language Model SFT Arithmetic Reasoning
START [202] Sampling Language Model SFT & RL Arithmetic Reasoning, Code Generation
@@
EffiLearner [206] Independent of Training Data Language Model External Tools Code Generation
@@
ProgCo [213] Independent of Training Data Language Model Program-driven Verification Instruction-following, Arithmetic Reasoning
@@ 3.2.3 Self Improvement
Reasoning LLMs exemplify a progression from weak to strong supervision, while traditional CoT fine-tuning faces challenges in scaling effectively. Self improvement, using the model’s exploration capabilities for self-supervision, gradually enhances LLMs performance [201] in tasks such as translation [189], mathematics [186, 190], and multimodal perception [193]. This approach fosters exploration and application within reasoning LLMs [214, 215, 216, 143]. A summary of Self Improvement method is presented in Table III.
@@ 3.2.5 Reinforcement Fine-Tuning
The success of R1 in unimodal reasoning has spurred its adaptation to MLLM, achieving SOTA performance across diverse domains (math [246, 247], medical imaging [248], segmentation [249]). Key strategies include rule-based reward systems to incentivize structured reasoning (e.g., step-wise validity/accuracy rewards in R1-VL [250]), modality-agnostic training frameworks (e.g., Vision-R1’s [251] cold-start CoT data generation), and efficient parameter utilization (e.g., MMR1 [246]’s 7B models rivaling larger proprietary counterparts). The work emphasizes transparency through explicit reasoning paths (e.g., MedVLM-R1’s [248] interpretable medical analysis) and open-source contributions (code [252], data [247], benchmarks [253, 254]), fostering reproducibility and MLLM-community improvement.
@@
Category Methods Model Attribute Incentivize Attribute Application & Benchmark
@@
LLM DeepSeek-R1-Zero[31] DeepSeek-V3 TT Rule-Outcome-Reward GPRO RL 800K Multiple Tasks
@@
Kimi k1.5[255] II TT Rule-Outcome-Reward $\text{PPO}^{*}$ RL & SFT Multiple Tasks
@@
RLAIF[270] PaLM 2 Extra-Small TT Rule-Outcome-Reward RLAIF RL Summary and Conversation Generation
MLLM RLHF-V[271] OmniLMM-12B II TT Process-Reward DDPO RL 1.4K Multiple Tasks
@@
R1V[274] Qwen2-VL,Qwen2.5-VL II TT Rule-Outcome-Reward GRPO RL 70K/70K/8K Multiple Tasks
VLM-R1[275] Qwen2.5-VL II TT Rule-Outcome-Reward GRPO RL 120K Multiple Tasks
LMM-R1[276] Qwen2.5-VL II TT Rule-Outcome-Reward PPO/RLOO RL 8K Multiple Tasks
Open-R1-Video[277] Qwen2-VL-7B II TT VV Rule-Outcome-Reward GRPO RL 4K Multiple Tasks
Easy-R1 [252] Qwen2.5-VL II TT Rule-Outcome-Reward GRPO RL 3K Multiple Tasks
Efficient-R1-VLLM[278] DeepSeek-VL2 MoE II TT Rule-Outcome-Reward GRPO RL Multiple Tasks
MMR1 [246] Qwen2.5-VL-7B II TT Rule-Outcome-Reward GRPO RL 6K Multiple Math Tasks
MedVLM-R1[248] Qwen2-VL-2B II TT Rule-Outcome-Reward GRPO RL 17.3K Medical VQA
MM-EUREKA [247] InternVL2.5-8B/38B II TT Rule-Outcome-Reward RLOO RL 54K Multiple Math Tasks
R1-Omni[279] HumanOmni-0.5B II TT VV Rule-Outcome-Reward GRPO RL & SFT 0.58K/15.3K Multimodal Emotion Recognition
R1-Onevision [253] Qwen2.5-VL-3B/7B II TT Rule-Outcome-Reward GRPO RL & SFT 155K Multiple Tasks
R1-VL[250] Qwen2-VL-2B/7B II TT Rule-Outcome-Reward stepGRPO RL & SFT 260K Multiple Tasks
VisualThinker R1 Zero [280] Qwen2-VL-2B, Qwen-2-VL-2B-Instruct II TT Rule-Outcome-Reward GRPO RL 218K CVBench
LALM AQA[281] Qwen2-Audio-7B-Instruc II TT VV Rule-Outcome-Reward GRPO RL 38k MMAU
Seg-Zero [249] Qwen2.5-VL-3B II TT Rule-Outcome-Reward GRPO RL 9K RefCOCO, ReasonSeg
Skywork R1V[282] InternViT-6B, DeepSeek-R1-Distill-Qwen-32B II TT Rule-Outcome-Reward GRPO RL & SFT Multiple Tasks
TimeZero [283] Qwen2.5-VL-7B II TT VV Rule-Outcome-Reward GRPO RL 37.4K Charades-STA, ActivityNet
Vision-R1[251] Qwen-2.5-VL-7B-Instruct II TT Rule-Outcome-Reward GRPO RL & SFT 10K/200K Multiple Tasks
Visual-RFT [284] Qwen2-VL-2B/7B II TT Rule-Outcome-Reward GRPO RL Visual Perception Tasks
Reason-RFT[285] Qwen-2-VL-2B/7B II TT Rule-Outcome-Reward GRPO RL Multiple Tasks
SEED-Bench-R1 [284] Qwen2-VL-Instruct-7B II TT VV Rule-Outcome-Reward GRPO RL 50.2K Video Understanding
@@
LLM Demystify-LongCoT[286] Llama-3.1-8B, Qwen2.5 -7B-Math TT Rule-Outcome-Reward PPO/Reinforce++ RL & SFT 7.5K Multiple Math, MMLU
@@
MD-CoT [288]
GraidientUnified [289]
Despite the strengths of RFT, it still faces the following challenges:
@@
    Unstable Long-CoT Generation: Long reasoning chains generated by RFT are prone to instability, including context overflow, failure to return final answers, and sensitivity to reward shaping [124]. For instance, methods like [286] inadvertently introduce cosine reward functions, which degrade performance with increased iterations. O1-Prune [292] uses post-hoc length pruning techniques [255] (via RL/SFT) to stabilize outputs. Recent VL models such as Open-R1-Video [277] and Seg-Zero [249] show instability in chain generation when handling long videos or fine-grained segmentation reasoning.
@@
    Theoretical and Empirical Analysis: It is essential to clarify the relationship between RL training and the capabilities of the base model. For instance, it should be determined whether emergent properties (e.g., Long-CoT) arise from RL optimization or are latent traits of the base model. However, this phenomenon of Long-CoT is often difficult to appear in multimodal situations (e.g., [274, 276, 252, 294]). Systematic studies on reward design principles (e.g., sparse vs. dense rewards, multi-objective balancing) should be conducted to avoid unintended behaviors such as reward hacking.